# Asynchronous Federated Reinforcement Learning with Policy Gradient Updates: Algorithm Design and Convergence Analysis

### v1

@@ A.1. Technical Lemmas
(27) $\|\sum_{i=1}^{n}\mathbf{a}_{i}\|^{2}~{}\leq~{}n\sum_{i=1}^{n}\|\mathbf{a}_{i}\|^{2}.$
Lemma A.2.
@@
(28) $1-\alpha_{k+1}~{}\leq~{}\frac{\alpha_{k+1}}{\alpha_{k}}.$
@@
(30) $\prod_{i=k}^{K-1}(1-\alpha_{i+1})~{}\leq$ $~{}\frac{\alpha_{K}}{\alpha_{k}}.$
Lemma A.3.
@@
(31) $\sum_{k=0}^{K-1}\eta_{k}\prod_{i=k+1}^{K-1}(1-\alpha_{i})~{}\leq~{}c(p,q)\frac{\eta_{K}}{\alpha_{K}},$
where $c(p,q)\coloneqq\frac{2^{q-p}}{1-p}a\exp{\big{(}(1-p)2^{p}a^{1-p}\big{)}}$ is a constant, and $a=\max\big{(}(\frac{q}{(1-p)2^{p}})^{\frac{1}{1-p}},(\frac{2(q-p)}{(1-p)^{2}})^{\frac{1}{1-p}}\big{)}$.
A.2. Proof of Theorem 4.3
Denote $\xi_{k}\coloneqq g(\widetilde{\tau}_{k},\widetilde{\theta}_{k})-\nabla J(\widetilde{\theta}_{k})$. First, according to the updating rules in Algo. 1 and Algo. 2 , we expand the error term in Lemma 4.6 as follows
(32) $e_{k}~{}=$ $~{}(1-\alpha_{k-\delta_{k}})d_{\delta_{k-1}}+(1-\alpha_{k-\delta_{k}})\big{(}\nabla J(\theta_{k-1})-\nabla J(\theta_{k})\big{)}+\alpha_{k-\delta_{k}}\big{(}g(\widetilde{\tau}_{k-\delta_{k}},\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\theta_{k})\big{)}$
@@
$+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\widetilde{\theta}_{k})\big{)}+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})\big{)}$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}-(1-\alpha_{k-\delta_{k}})\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})-\alpha_{k-\delta_{k}}\nabla^{2}J(\theta_{k})(\widetilde{\theta}_{k}-\theta_{k})}$
${{=}}$ $~{}\alpha_{k-\delta_{k}}\xi_{k-\delta_{k}}+(1-\alpha_{k-\delta_{k}})e_{k-1}+(1-\alpha_{k-\delta_{k}})\big{(}\nabla J(\theta_{k-1})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})\big{)}$
$+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\widetilde{\theta}_{k})\big{)}+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})\big{)}.$
Remark: The Hessian correction terms (in blue) are equal to 00 according to our updating rules (delay-adaptive lookahead update) in Step 2 of Algo. 2. We design this technique to cancel out the second-order terms, and thus achieve the desired convergence rate.
Next, we denote
(33) $A_{k}~{}\coloneqq$ $~{}\nabla J(\theta_{k-1})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k}),$
@@
With the smoothness in Lemma 4.4 and the non-increasing learning rates, we have
(34) $\|A_{k}\|~{}$ $\leq~{}L_{h}\|\theta_{k-1}-\theta_{k}\|^{2}=L_{h}\eta_{k-1}^{2},$
$\|B_{k}\|~{}$ $\leq~{}L_{h}\|\widetilde{\theta}_{k}-\theta_{k}\|^{2}=L_{h}\frac{(1-\alpha_{k-\delta_{k}})^{2}}{\alpha_{k-\delta_{k}}^{2}}\eta_{k-1}^{2},$
$\|C_{k}\|~{}$ $\leq~{}L_{g}\|\widetilde{\theta}_{k-\delta_{k}}-\widetilde{\theta}_{k}\|$
@@
$\leq~{}L_{g}\sum_{i=k-\delta_{k}}^{k-1}(\frac{1}{\alpha_{i+1-\delta_{i+1}}}\eta_{i}+\frac{1-\alpha_{i-\delta_{i}}}{\alpha_{i-\delta_{i}}}\eta_{i-1})$
@@
$=~{}\delta_{k}L_{g}\frac{2\eta_{k-\delta_{k}}}{\alpha_{k-1}}.$
Denote $\beta_{k}\coloneqq\prod_{i=k+1}^{K}(1-\alpha_{i-\delta_{i}})$ with $\beta_{K}=1$. Unrolling the recursion (32), we have
(35) $e_{K}~{}$ $=~{}\beta_{0}e_{0}+\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\xi_{k-\delta_{k}}+\sum_{k=1}^{K}(1-\alpha_{k-\delta_{k}})\beta_{k}A_{k}+\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}B_{k}.$
@@
(36) $\mathbb{E}[\|e_{K}\|]~{}\leq$ $~{}\beta_{0}\mathbb{E}[\|e_{0}\|]+\mathbb{E}\left[\left\|\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\xi_{k-\delta_{k}}\right\|^{2}\right]^{\frac{1}{2}}+\sum_{k=1}^{K}(1-\alpha_{k-\delta_{k}})\beta_{k}\mathbb{E}[\|A_{k}\|]$
@@
(${{\leq}}$ $~{}\beta_{0}\sigma_{g}+\left(\sum_{k=1}^{K}\alpha_{k-\delta_{k}}^{2}\beta_{k}^{2}\mathbb{E}\left[\left\|\xi_{k-\delta_{k}}\right\|^{2}\right]\right)^{\frac{1}{2}}+L_{h}\sum_{k=1}^{K}(1-\alpha_{k-\delta_{k}})\beta_{k}\eta_{k-1}^{2}$
$+L_{h}\sum_{k=1}^{K}\beta_{k}\frac{(1-\alpha_{k-\delta_{k}})^{2}}{\alpha_{k-\delta_{k}}}\eta_{k-1}^{2}+2L_{g}\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\delta_{k}\frac{\eta_{k-\delta_{k}}}{\alpha_{k-1}}$
@@
(${{\leq}}$ $~{}\alpha_{K}\sigma_{g}+\left(\sum_{k=1}^{K}\alpha_{k-\delta_{k}}^{2}\beta_{k}^{2}\right)^{\frac{1}{2}}\sigma_{g}+\frac{9L_{h}}{4}\sum_{k=1}^{K}\frac{\eta_{k}^{2}}{\alpha_{k}}\beta_{k}+2L_{g}\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\delta_{k}\frac{\eta_{k-\delta_{k}}}{\alpha_{k-1}}$
(${{\leq}}$ $~{}\alpha_{K}\sigma_{g}+c_{1}\sqrt{\alpha_{K}}\sigma_{g}+\frac{9}{4}c_{2}L_{h}\frac{\eta_{K}^{2}}{\alpha_{K}^{2}}+2c_{3}L_{g}\frac{\eta_{K}^{2}}{\alpha_{K}}\bar{\delta},$
where $c_{1}\coloneqq\sqrt{c(\frac{4}{5},\frac{8}{5})}$, $c_{2}\coloneqq c(\frac{4}{5},\frac{6}{5})$, $c_{3}\coloneqq c(\frac{4}{5},0)$ are constants, and the values of $c(\cdot,\cdot)$ are defined in Lemma A.3. After plugging the result into Lemma 4.6 and using Lemma 4.5, we achieve
(37) $J^{\star}-\mathbb{E}[J(\theta_{k+1})]~{}\leq$ $~{}(1-\frac{\sqrt{2\mu}\eta_{k}}{3})\big{(}J^{\star}-\mathbb{E}[J(\theta_{k})]\big{)}+\frac{\eta_{k}}{3}\epsilon_{g}+\frac{8}{3}\eta_{k}\mathbb{E}[\|e_{k}\|]+\frac{L_{g}}{2}\eta_{k}^{2}$
(${{\leq}}$ $~{}(1-\frac{\sqrt{2\mu}\eta_{k}}{3})\big{(}J^{\star}-\mathbb{E}[J(\theta_{k})]\big{)}+\frac{\eta_{k}}{3}\epsilon_{g}+\frac{L_{g}}{2}\eta_{k}^{2}$
$+\frac{8}{3}\eta_{k}\big{(}\alpha_{k}\sigma_{g}+c_{1}\sqrt{\alpha_{k}}\sigma_{g}+\frac{9}{4}c_{2}L_{h}\frac{\eta_{k}^{2}}{\alpha_{k}^{2}}+2c_{3}L_{g}\frac{\eta_{k}^{2}}{\alpha_{k}}\bar{\delta}\big{)}.$
@@
(38) $J^{\star}-\mathbb{E}[J(\theta_{K})]~{}\leq$ $~{}\frac{\eta_{0}}{3}\epsilon_{g}+\frac{J^{\star}-\mathbb{E}[J(\theta_{0})]}{(K+1)^{2}}+c_{3}\frac{16\eta_{0}^{2}}{3}\frac{L_{g}}{(K+1)^{\frac{7}{5}}}\bar{\delta}+\frac{\eta_{0}^{2}}{2}\frac{L_{g}}{K+1}+\frac{8\eta_{0}}{3}\frac{\sigma_{g}}{(K+1)^{\frac{4}{5}}}$
$+c_{1}\frac{8\eta_{0}}{3}\frac{\sigma_{g}}{(K+1)^{\frac{2}{5}}}+6c_{2}\eta_{0}^{3}\frac{L_{h}}{(K+1)^{\frac{2}{5}}}.$

### v5

@@ B.2 Technical Lemmas
$\|\sum_{i=1}^{n}\mathbf{a}_{i}\|^{2}~{}\leq~{}n\sum_{i=1}^{n}\|\mathbf{a}_{i}\|^{2}.$ (15)
Lemma B.5.
@@
$1-\alpha_{k+1}~{}\leq~{}\frac{\alpha_{k+1}}{\alpha_{k}}.$ (16)
@@
$\prod_{i=k}^{K-1}(1-\alpha_{i+1})~{}\leq$ $~{}\frac{\alpha_{K}}{\alpha_{k}}.$ (18)
Lemma B.6.
@@
$\sum_{k=0}^{K-1}\eta_{k}\prod_{i=k+1}^{K-1}(1-\alpha_{i})~{}\leq~{}c(p,q)\frac{\eta_{K}}{\alpha_{K}},$ (19)
where $c(p,q)\coloneqq\frac{2^{q-p}}{1-p}a\exp{\big{(}(1-p)2^{p}a^{1-p}\big{)}}$ is a constant with specified $p$ and $q$, and $a=\max\big{(}(\frac{q}{(1-p)2^{p}})^{\frac{1}{1-p}},(\frac{2(q-p)}{(1-p)^{2}})^{\frac{1}{1-p}}\big{)}$.
B.3 Key Lemmas
In this subsection, we list four useful (key) lemmas to construct the proofs of Theorem 5.1 and Theorem 5.2.
Under Assumption B.1 on score functions, the following lemma holds based on the results (Lemma 5.4) in Zhang et al. (2020).
Lemma B.7.
$\|\nabla J(\theta)-\nabla J(\theta^{\prime})\|~{}\leq$ $~{}L_{g}\|\theta-\theta^{\prime}\|,$ (20)
Under Assumption B.1, B.2 and B.3, we utilize the result (Lemma 4.7) in Ding et al. (2022) as the Lemma B.8.
Lemma B.8.
(Relaxed weak gradient domination) Under Assumptions B.1, B.2 and B.3, it holds that
$\|\nabla J(\theta)\|+\epsilon_{g}~{}\geq~{}\sqrt{2\mu}(J^{\star}-J(\theta)),$ (21)
where $\epsilon_{g}=\frac{\mu_{F}\sqrt{\epsilon_{{\rm bias}}}}{M_{g}(1-\gamma)}$ and $\mu=\frac{\mu_{F}^{2}}{2M_{g}^{2}}$.
Based on Lemma B.7, we derive our milestone (new), the ascent lemma with delayed updates, as follows:
Lemma B.9.
(Ascent Lemma with Delay) Under Assumptions B.1, it holds that
$-J(\theta_{k+1})~{}\leq~{}-J(\theta_{k})-\frac{1}{3}\eta_{k}\|\nabla J(\theta_{k})\|+\frac{8}{3}\eta_{k}\|e_{k}\|+\frac{L_{g}}{2}\eta_{k}^{2},$ (22)
where $e_{k}\coloneqq d_{k-\delta_{k}}-\nabla J(\theta_{k})$.
Proof.
$-J(\theta_{k+1}){}$ $\leq~{}-J(\theta_{k})-\langle\nabla J(\theta_{k}),\theta_{k+1}-\theta_{k}\rangle+\frac{L_{g}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$ (23)
If $\|e_{k}\|\leq\frac{1}{2}\|\nabla J(\theta_{k})\|$, we have
$-\frac{\langle\nabla J(\theta_{k}),d_{k-\delta_{k}}\rangle}{\|d_{k-\delta_{k}}\|}{}$ $=~{}-\frac{\|\nabla J(\theta_{k})\|^{2}+\langle\nabla J(\theta_{k}),e_{k}\rangle}{\|d_{k-\delta_{k}}\|}$ (24)
$\leq~{}\frac{-\|\nabla J(\theta_{k})\|^{2}+\frac{1}{2}\|\nabla J(\theta_{k})\|^{2}}{\|\nabla J(\theta_{k})\|+\|e_{k}\|}$
$\leq~{}-\frac{1}{3}\|\nabla J(\theta_{k})\|.$
If $\|e_{k}\|\geq\frac{1}{2}\|\nabla J(\theta_{k})\|$, we have
$-\frac{\langle\nabla J(\theta_{k}),d_{k-\delta_{k}}\rangle}{\|d_{k-\delta_{k}}\|}{}$ $\leq~{}\|\nabla J(\theta_{k})\|$ (25)
$=~{}-\frac{1}{3}\|\nabla J(\theta_{k})\|+\frac{4}{3}\|\nabla J(\theta_{k})\|$
$\leq~{}-\frac{1}{3}\|\nabla J(\theta_{k})\|+\frac{8}{3}\|e_{k}\|.$
Combining these two conditions, and plugging the result into equation 23, the lemma can be proved as follows
$-J(\theta_{k+1}){}$ $\leq~{}-J(\theta_{k})-\eta_{k}\frac{\langle\nabla J(\theta_{k}),d_{k-\delta_{k}}\rangle}{\|d_{k-\delta_{k}}\|}+\frac{L_{g}}{2}\eta_{k}^{2}$ (26)
Next, we construct the relationship between the average concurrency $\bar{\omega}$ and the average delay $\bar{\delta}$ in Lemma B.10. This lemma gives the boundary (our result) of delays as a corollary of the result in Koloskova et al. (2022).
Lemma B.10.
The average delay $\bar{\delta}$ depends on the average concurrency $\bar{\omega}$, and they can be upper bounded as
$\bar{\delta}~{}=~{}\frac{K+1}{K-1+|\mathcal{C}_{K}|}\bar{\omega}~{}\leq~{}\bar{\omega}~{}\leq~{}N.$ (27)
Proof.
$\sum_{i=0}^{k}\delta_{i}+\sum_{i\in\mathcal{C}_{k+1}\setminus\{j_{k+1}\}}\delta^{i}_{k+1}{}$ $=~{}\sum_{i=0}^{k-1}\delta_{i}+\sum_{i\in\mathcal{C}_{k}\setminus\{j_{k}\}}\delta^{i}_{k}+\omega_{k+1}.$ (28)
We note that there is no delay at the initial step (00-th iteration) of the algorithm. Therefore, we have $\delta^{i}_{0}=0$ for all agents. Unrolling the above expression, at the $K$-th step, we have
$\sum_{i=0}^{K-1}\delta_{i}+\sum_{i\in\mathcal{C}_{K}\setminus\{j_{K}\}}\delta^{i}_{K}{}$ $=~{}\sum_{i=0}^{K}\omega_{k+1}~{}=~{}(K+1)\bar{\omega}.$ (29)
According to equation 7 and $|\mathcal{C}_{K}|\geq 2$, we achieve
$\bar{\delta}~{}=~{}\frac{K+1}{K-1+|\mathcal{C}_{K}|}\bar{\omega}~{}\leq~{}\bar{\omega}~{}\leq~{}N.$ (30)
In practice, we need to use all the resources to speed up the training process. Thus, all agents engage in training, and $\bar{\omega}=\omega_{\max}=N$. Since the maximum concurrency is equal to the number of agents $N$, we have the upper boundary of the average delay as $\bar{\delta}\leq\omega_{\max}\leq N$.
Notably, we do not make any assumption on the largest delay $\delta_{\max}$, while we achieve the upper boundary of the average delay $\bar{\delta}$.
B.4 Proof of Theorem 5.1 (Global Convergence Rate)
Under Assumption B.1 and Assumption B.2, we derive the global convergence rate of the proposed AFedPG.
First, we denote the difference between the policy gradient estimation $g(\widetilde{\tau}_{k},\widetilde{\theta}_{k})$ and the true policy gradient as
$\xi_{k}\coloneqq g(\widetilde{\tau}_{k},\widetilde{\theta}_{k})-\nabla J(\widetilde{\theta}_{k}),$ (31)
and the expectation of the norm is bounded by $\sigma_{g}$.
Second, according to the updating rules in Algorithm 1 and Algorithm 2 , we expand the error term in Lemma B.9 as follows
$e_{k}~{}=$ $~{}(1-\alpha_{k-\delta_{k}})d_{k-1-\delta_{k-1}}-\nabla J(\theta_{k})+\alpha_{k-\delta_{k}}g(\widetilde{\tau}_{k-\delta_{k}},\widetilde{\theta}_{k-\delta_{k}})$ (32)
@@ B.3 Key Lemmas
$+\alpha_{k-\delta_{k}}\big{(}g(\widetilde{\tau}_{k-\delta_{k}},\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\theta_{k})\big{)}$
[…]
$+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\theta_{k})\big{)}$
@@
$+(1-\alpha_{k-\delta_{k}})\big{(}\nabla J(\theta_{k-1})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})\big{)}$
$+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})\big{)}$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-(1-\alpha_{k-\delta_{k}})\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})-\alpha_{k-\delta_{k}}\nabla^{2}J(\theta_{k})(\widetilde{\theta}_{k}-\theta_{k})}$
${{=}}$ $~{}\alpha_{k-\delta_{k}}\xi_{k-\delta_{k}}+(1-\alpha_{k-\delta_{k}})e_{k-1}+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\widetilde{\theta}_{k})\big{)}$
$+(1-\alpha_{k-\delta_{k}})\big{(}\nabla J(\theta_{k-1})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})\big{)}$
$+\alpha_{k-\delta_{k}}\big{(}\nabla J(\widetilde{\theta}_{k})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})\big{)}.$
Thus, the error term $e_{k}$ can be written in a recursive way (contains $e_{k-1}$) with serval terms that contain policy gradients. We aim to derive the upper boundary for each term at the next step.
Remark: The Hessian correction terms (in blue) are equal to 00 according to our updating rules (delay-adaptive lookahead update) in Step 8 of Algorithm 1. We design this technique to cancel out the second-order terms, and thus achieve the desired convergence rate. Without our technique, it would be hard to bound the above errors.
Next, we denote each term in equation 32 separately as follows
$A_{k}~{}\coloneqq$ $~{}\nabla J(\theta_{k-1})-\nabla J(\theta_{k})+\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k}),$ (33)
@@
Now, we start to bound each term. With the smoothness of the expected discounted return function in Lemma B.7 and the non-increasing learning rates, we have
$\|A_{k}\|{}$ $\leq~{}L_{h}\|\theta_{k-1}-\theta_{k}\|^{2}=L_{h}\eta_{k-1}^{2},$ (34)
$\|B_{k}\|{}$ $\leq~{}L_{h}\|\widetilde{\theta}_{k}-\theta_{k}\|^{2}=L_{h}\frac{(1-\alpha_{k-\delta_{k}})^{2}}{\alpha_{k-\delta_{k}}^{2}}\eta_{k-1}^{2},$
$\|C_{k}\|{}$ $\leq~{}L_{g}\|\widetilde{\theta}_{k-\delta_{k}}-\widetilde{\theta}_{k}\|$
@@
$=~{}L_{g}\sum_{i=k-\delta_{k}}^{k-1}\left\|\frac{1}{\alpha_{i+1-\delta_{i+1}}}({\theta}_{i+1}-{\theta}_{i})+\frac{1-\alpha_{i-\delta_{i}}}{\alpha_{i-\delta_{i}}}({\theta}_{i}-{\theta}_{i-1})\right\|$
$\leq~{}L_{g}\sum_{i=k-\delta_{k}}^{k-1}\left\|\frac{1}{\alpha_{i+1-\delta_{i+1}}}({\theta}_{i+1}-{\theta}_{i})\right\|+\left\|\frac{1-\alpha_{i-\delta_{i}}}{\alpha_{i-\delta_{i}}}({\theta}_{i}-{\theta}_{i-1})\right\|$
$=~{}L_{g}\sum_{i=k-\delta_{k}}^{k-1}(\frac{1}{\alpha_{i+1-\delta_{i+1}}}\eta_{i}+\frac{1-\alpha_{i-\delta_{i}}}{\alpha_{i-\delta_{i}}}\eta_{i-1})$
@@
$=~{}\delta_{k}L_{g}\frac{2\eta_{k-\delta_{k}}}{\alpha_{k-1}},$
where the equalities are simple plugins according to the updating rules in Algorithm 1. The last step in equation 34 happens, because there is no index $i$ inside the summation operation, and it sums a constant for $\delta_{k}$ times.
We denote $\beta_{k}\coloneqq\prod_{i=k+1}^{K}(1-\alpha_{i-\delta_{i}})$ with $\beta_{K}=1$. Next, unrolling the recursion equation 32, we have the error at the $K$-th step as follows
$e_{K}{}$ $=~{}\beta_{0}e_{0}+\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\xi_{k-\delta_{k}}+\sum_{k=1}^{K}(1-\alpha_{k-\delta_{k}})\beta_{k}A_{k}+\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}(B_{k}+C_{k}).$ (35)
@@
$\mathbb{E}[\|e_{K}\|]~{}\leq$ $~{}\beta_{0}\mathbb{E}[\|e_{0}\|]+\left(\mathbb{E}\left[\left\|\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\xi_{k-\delta_{k}}\right\|^{2}\right]\right)^{\frac{1}{2}}+\sum_{k=1}^{K}(1-\alpha_{k-\delta_{k}})\beta_{k}\mathbb{E}[\|A_{k}\|]$ (36)
@@
${{\leq}}$ $~{}\beta_{0}\sigma_{g}+\left(\sum_{k=1}^{K}\alpha_{k-\delta_{k}}^{2}\beta_{k}^{2}\mathbb{E}\left[\left\|\xi_{k-\delta_{k}}\right\|^{2}\right]\right)^{\frac{1}{2}}+L_{h}\sum_{k=1}^{K}(1-\alpha_{k-\delta_{k}})\beta_{k}\eta_{k-1}^{2}$
$+L_{h}\sum_{k=1}^{K}\beta_{k}(1-\alpha_{k-\delta_{k}})^{2}\frac{\eta_{k-1}^{2}}{\alpha_{k-\delta_{k}}}+2L_{g}\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\delta_{k}\frac{\eta_{k-\delta_{k}}^{2}}{\alpha_{k-1}}$
[…]
$+2L_{g}\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\delta_{k}\frac{\eta_{k-\delta_{k}}^{2}}{\alpha_{k-1}}$
[…]
where $c_{1}\coloneqq 2\sqrt{c(\frac{4}{5},\frac{4}{5})}$, $c_{2}\coloneqq c(\frac{4}{5},\frac{6}{5})$, $c_{3}\coloneqq 8\sqrt{c(\frac{4}{5},\frac{16}{5})}$ are constants, and the values of $c(\cdot,\cdot)$ are defined in Lemma B.6.
Notably, as we state in the last paragraph in Section 5, $t_{i}$ is pre-determined by the computation resource at agent $i$ and is fixed. Thus, the delay $\delta_{k}$ is pre-determined by the system. It is not a random variable, but unknown until given the exact system setting. This allows us to derive the first inequality in equation 36. This pre-determined property is also suitable for all the derivations below.
We explain the boundary derivation details here. We first derive the first term of the boundary in equation 36 as follows
$\beta_{0}~{}=$ $~{}\prod_{i=1}^{K}(1-\alpha_{i-\delta_{i}})$ (37)
[…]
${{\leq}}$ $~{}\frac{\alpha_{K}}{\alpha_{0}}$
[…]
In the training process, at step $k$, when an agent sends the update to the server, as long as the agent communicates with the server during the last half training process, the delay $\delta_{k}\leq\frac{k}{2}$. This becomes almost certain when $k$ is large and $k$ is usually much larger than the upper bound of the currency $N$. Thus, we have
$\alpha_{k-\delta_{k}}~{}=$ $~{}(\frac{1}{k+1-\delta_{k}})^{\frac{4}{5}}$ (38)
[…]
We then derive the second term of the boundary in equation 36 as follows
[…]
${{\leq}}$ $~{}4\sum_{k=1}^{K}\alpha_{k}^{2}\prod_{i=k+1}^{K-1}(1-\alpha_{i})\frac{\alpha_{K}}{\alpha_{k}}$
[…]
${{\leq}}$ $~{}4\alpha_{K}c(\frac{4}{5},\frac{4}{5})\frac{\alpha_{K}}{\alpha_{K}}$
[…]
Next, we derive the third term of the boundary in equation 36 as follows
$\sum_{k=1}^{K}\frac{\eta_{k}^{2}}{\alpha_{k}}\beta_{k}~{}=$ $~{}\eta_{0}^{2}\sum_{k=1}^{K}(\frac{1}{k+1})^{\frac{6}{5}}\beta_{k}$ (40)
[…]
${{\leq}}$ $~{}\eta_{0}^{2}c(\frac{4}{5},\frac{6}{5})(\frac{1}{K+1})^{\frac{2}{5}}$
[…]
At last, we derive the fourth term of the boundary in equation 36. We use Cauchy–Schwarz inequality and the fact that the second norm of a vector is always equal or smaller than the first norm.
$\sum_{k=1}^{K}\alpha_{k-\delta_{k}}\beta_{k}\delta_{k}\frac{\eta_{k-\delta_{k}}^{2}}{\alpha_{k-1}}~{}\leq$ $~{}\sum_{k=1}^{K}\eta_{k-\delta_{k}}^{2}\beta_{k}\delta_{k}$
[…]
${{\leq}}$ $~{}4\eta_{0}^{2}K\bar{\delta}\left(\alpha_{K}c(\frac{4}{5},\frac{16}{5})(\frac{1}{K+1})^{\frac{12}{5}}\right)^{\frac{1}{2}}$
[…]
Now, we derive the final convergence rate. After plugging the result into Lemma B.9 and using Lemma B.8 (Ascent Lemma with Delay), we achieve the following inequality
$J^{\star}-\mathbb{E}[J(\theta_{k+1})]~{}\leq$ $~{}(1-\frac{\sqrt{2\mu}\eta_{k}}{3})\big{(}J^{\star}-\mathbb{E}[J(\theta_{k})]\big{)}+\frac{\eta_{k}}{3}\epsilon_{g}+\frac{8}{3}\eta_{k}\mathbb{E}[\|e_{k}\|]+\frac{L_{g}}{2}\eta_{k}^{2}$ (42)
${{\leq}}$ $~{}(1-\frac{\sqrt{2\mu}\eta_{k}}{3})\big{(}J^{\star}-\mathbb{E}[J(\theta_{k})]\big{)}+\frac{\eta_{k}}{3}\epsilon_{g}+\frac{L_{g}}{2}\eta_{k}^{2}$
$+\frac{8}{3}\eta_{k}\big{(}\alpha_{k}\sigma_{g}+c_{1}\sqrt{\alpha_{k}}\sigma_{g}+\frac{9}{4}c_{2}L_{h}\frac{\eta_{k}^{2}}{\alpha_{k}^{2}}+c_{3}L_{g}\frac{\eta_{k}^{2}}{\alpha_{k}^{1.75}}\bar{\delta}\big{)}.$
@@
$J^{\star}-\mathbb{E}[J(\theta_{K})]~{}\leq$ $~{}\frac{\eta_{0}}{3}\epsilon_{g}+\frac{J^{\star}-\mathbb{E}[J(\theta_{0})]}{(K+1)^{2}}+c_{3}\frac{8\eta_{0}^{3}}{3}\frac{L_{g}}{(K+1)^{\frac{3}{5}}}\bar{\delta}+\frac{\eta_{0}^{2}}{2}\frac{L_{g}}{K+1}$ (43)
$+\frac{8\eta_{0}}{3}\frac{\sigma_{g}}{(K+1)^{\frac{4}{5}}}+c_{1}\frac{8\eta_{0}}{3}\frac{\sigma_{g}}{(K+1)^{\frac{2}{5}}}+6c_{2}\eta_{0}^{3}\frac{L_{h}}{(K+1)^{\frac{2}{5}}}.$